Let V be a Banach space and let M ⊆ V be a proper subspace of V (i.e. M not equal to V ). Prove that if v ∈ V and v /∈ M then there is a φ ∈ V* such that φ(v) = 1 and φ(w) = 0 for every w ∈ M.


1
Expert's answer
2021-02-24T06:10:16-0500

First of all, we need to have "\bar M \neq V, v\notin \bar M", otherwise it would not be possible to construct such a form.

Define "N=M\oplus \mathbb{K}v" and now for any "w \in N, w=\lambda_M\cdot m+\lambda_v\cdot v", where "m\in M". Now we can define a linear form "\phi_N :(\lambda_M \cdot m + \lambda_v \cdot v)\mapsto \lambda_v". Let's prove that it is continuous by calculating it's norm "||\phi_N|| = \sup_{w\in N}\frac{|\lambda_v |}{||\lambda_M\cdot m + \lambda_v \cdot v||}", "||\phi_N|| =\sup_{w\in N} \frac{1}{||v+\frac{\lambda_M \cdot m}{|\lambda_v|}||}" , the vector "\frac{\lambda_M}{|\lambda_v|}m" is in "M" and as "v\notin \bar M", the expression below is bounded below by some constant (as otherwise the vector "v" could be approached by "-\frac{\lambda_M}{|\lambda_v|}m" as closely, as we want and thus "v\in \bar M"). Therefore the norm "||\phi_N||" is bounded, so "\phi_N" is continuous. Now by Hahn-Banach's theorem we can extend "\phi_N" on "V" such that the extension "\phi|_M=0, \phi(v)=1, \phi" is continuous.